OpenStudy (anonymous):

the quantity of x plus 16 all over 3 = 3x

OpenStudy (anonymous):

@dpasingh please make sure to provide an explanation, no direct answers

OpenStudy (anonymous):

\[\frac{(x+16)}{3}=3x \rightarrow x+16 = 9x \rightarrow 16 = 9x -x \] i.e. 8x = 16 i.e. x= 16/8 i.e. x= 2